(1 point) Solve the system 3 9 da dt 2 -1 -3 2 with x(0) 4 Give your solution in real form. 21 = 22 =
Answer #1

3 a Solution: da dt (0) = [ -1 -3 Denote A [ Flest we find eigenvalues for A. det (A-XI) = det -3-] = (3-1)(-3-1) +9 :.det (APut S=0 1 X = -1, 242=O :. + O ot :. (t) = Cu ) +C teo e E] ] [i] £ ܕܝܚ - C cal + D] + 34+ (-3t-1) -34, +(z (-3t-1 1) 634 ci